# xTED: Cross-Domain Adaptation via Diffusion-Based Trajectory Editing

### v2

@@ Abstract
Reusing pre-collected data from different domains is an appealing solution for decision-making tasks that have insufficient data in the target domain but are relatively abundant in other related domains. Existing cross-domain policy transfer methods mostly aim at learning domain correspondences or corrections to facilitate policy learning, such as learning domain/task-specific discriminators, representations, or policies. This design philosophy often results in heavy model architectures or task/domain-specific modeling, lacking flexibility. This reality makes us wonder: can we directly bridge the domain gaps universally at the data level, instead of relying on complex downstream cross-domain policy transfer models? In this study, we propose the Cross-Domain Trajectory EDiting (xTED) framework that employs a specially designed diffusion model for cross-domain trajectory adaptation. Our proposed model architecture effectively captures the intricate dependencies among states, actions, and rewards, as well as the dynamics patterns within target data. By utilizing the pre-trained diffusion as a prior, source domain trajectories can be transformed to match with target domain properties while preserving original semantic information. This process implicitly corrects underlying domain gaps, enhancing state realism and dynamics reliability in the source data, and allowing flexible incorporation with various downstream policy learning methods. Despite its simplicity, xTED demonstrates superior performance in extensive simulation and real-robot experiments.
@@ 1 Introduction
Policies learned on insufficient data from a single domain easily end up with undesirable performance, especially in real systems (Guiochet et al., 2017; Zhan et al., 2022) where performing reinforcement/imitation learning (RL/IL) in the real world (i.e. target domain) yields laborious and costly data collection, notorious reset, and reward specification issues (Zhu et al., 2020). Faced with restricted data acquisition in reality, people resort to incorporating data generated with simulation or pre-collected from other domains (i.e. source domains) into policy learning for greater synergy (Niu et al., 2022; 2023; Open X-Embodiment et al., 2023). However, source domains invariably bear multiple domain gaps (Niu et al., 2024), such as appearance (Tobin et al., 2017) and dynamics gaps (Peng et al., 2018), as well as morphology gaps for embodied agents (Gupta et al., 2021). These domain discrepancies would reduce the availability of source data since directly augmenting target data with unprocessed source data sometimes leads to negative effects on policy learning.
@@
Existing cross-domain policy learning methods tend to design domain-specific policy transfer models equipped with domain correspondences (Zhang et al., 2020), corrections (Eysenbach et al., 2021), or discriminations (Stadie et al., 2016; Sharma et al., 2019) that either rely on task-specific architecture designs (e.g. special encoders (Mueller et al., 2018; Wang et al., 2022a) and domain-specific regularizations (Desai et al., 2020; Eysenbach et al., 2021; Niu et al., 2022; Xue et al., 2023)), or are only applicable to specific domain of data (e.g. adaptable vision encoder design only applicable to image inputs (Rao et al., 2020). Additionally, the domain-specific discriminators or mappings (Stadie et al., 2016; Liu et al., 2018) sometimes require tedious efforts to refresh for accomodating multiple source domains, which substantially hinders data reutilization efficiency and efficacy. These limitations highlight the need for a more generic and flexible approach that addresses domain gaps more directly.
Figure 1: The connections and distinctions between image editing and trajectory editing.
@@
Above inadequately addressed issues make us wonder: Rather than relying on sophisticated and domain-specific policy transfer models, is it possible to bridge domain gaps at the data level? If source trajectories can be transformed to minimize domain gaps and seamlessly utilized for policy learning, any policy learning method can be flexibly chosen with primary focus on task-relevant considerations, without the burden of cross-domain design complexities. To achieve this goal, we tend to correct the biased domain modeling within trajectories based on target data while preserving valuable and primitive information from source domains. With these pursuits, a natural analogy can be drawn that domain transfer at the trajectory level parallels style transfer in image editing (Meng et al., 2022), where only the visual style and aesthetic elements of an image are transformed without altering its narrative content, such as converting a stroke-painted castle into a photorealistic one. However, as depicted in Fig. 1, source-domain trajectories face not only biased observation styles but also altered viewpoints and inaccurate transition dynamics over time. Moreover, trajectories are composed of fundamentally heterogeneous elements—such as observations, actions, and rewards—that, if naïvely treated as equivalent pixels in an image, would obscure the recognition of their respective inherence and delicate interactive dependencies. Thus, despite sharing a similar high-level philosophy, this editing solution is not directly applicable to cross-domain policy transfer due to the intrinsic heterogeneity and dependencies within trajectories and the complex domain discrepancies.
@@
To address these distinctions, we propose the Cross-Domain Trajectory EDiting (xTED) paradigm with a novel diffusion model architecture tailored for decision-making data, which effectively captures the target trajectory distribution as a prior. The architecture illustrated in Fig. 2 employs a separate encoding-decoding approach to recognize states, actions, and rewards, each possessing fundamentally different physical meanings and unique internal consistency patterns. To effectively capture the dynamics patterns within trajectories, we implement dependency structure modeling mechanisms that capture the interactive relationships among states, actions, and rewards. With the assistance of above designs, we successfully extend the philosophy in diffusion-based image editing (Huang et al., 2024) to decision-making data, equipping xTED pipeline with three simple steps: (1) train diffusion model on the target data; (2) distort source data with noises and then denoise them with the pre-trained diffusion model; (3) incorporate edited source data into target data for policy learning with any algorithm at will. Through extensive experiments, we show that incorporating source data edited with xTED consistently yields performance improvements over training solely on target data while directly adding unprocessed source data often results in significant performance degradation, particularly in real-robot manipulation tasks.
@@ 2.1 Cross-Domain Policy Adaptation
Addressing the challenges posed by domain gaps has long been established as an essential task for the prolonged development of data-hungry policy learning methods and models (Niu et al., 2024). The most straightforward approach is to construct direct mappings for state and action space between source and target domains (Liu et al., 2018; Kim et al., 2020; Zhang et al., 2020; Raychaudhuri et al., 2021; Wang et al., 2022b). An alternative is to learn domain-agnostic task-relevant embedding with mutual information criterion (Franzmeyer et al., 2022) and explicit domain discrimination (Stadie et al., 2016; Sharma et al., 2019), which can also be achieved with temporal contrastive learning methods (Sermanet et al., 2018; Dwibedi et al., 2019; Yang et al., 2023; Choi et al., 2023; Li et al., 2024) to enable representation highly dependent on task progress while neglecting domain-variant information. Other than representation learning, it is also preferable to directly regularize the policy learning process in situations with mild observation gaps, by reward augmentation (Eysenbach et al., 2021; Liu et al., 2022; Xue et al., 2023) and reweighting the value update (Niu et al., 2022; 2023; Xu et al., 2023a). The complicated domain/task-specific designs hinder convenient reuse and fine-tuning of those models to accommodate data from multiple source domains available at different stages of training (Niu et al., 2024). In contrast, xTED introduces a generic and flexible approach that handles domain gaps at the data level, avoiding any task-specific design and domain-specific fine-tuning while applicable to multiple source domains of data and freeing the choices of any upstream observation encoder and downstream policy learning method.
@@ 2.2 Diffusion Models for Decision Making
Diffusion models have been applied across various decision-making settings, including generating multi-modal policies (Chi et al., 2023; Wang et al., 2023), single-step transitions (Lu et al., 2023), subgoals (Black et al., 2024), trajectories (Janner et al., 2022; Ajay et al., 2023a; He et al., 2023), and videos for planning (Ajay et al., 2023b; Yang et al., 2024; Du et al., 2023; 2024). However, prior works primarily focus on single-domain generation tasks, such as data augmentation, without addressing cross-domain challenges. In contrast, xTED introduces a distinct setting by employing diffusion models to adjust source trajectories so that they more closely resemble target domain characteristics while preserving useful primitive information from source domains. As demonstrated in Section 5.5, this setting is highly sample-efficient, as it does not require the extensive training data needed by diffusion-based generation tasks that aim to generate data from modeled target distribution.
@@ 3 Preliminaries
Notations.
For convenience, we introduce the relevant notations and elements of decision trajectories with the standard formulation of Markov Decision Process (MDP). A finite-horizon MDP is defined as a tuple $\mathcal{M}:=\langle\mathcal{S},\mathcal{A},T,R,H,\gamma\rangle$ where $\mathcal{S}$ and $\mathcal{A}$ are state and action spaces, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition dynamics; $R:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathbb{R}}$ is the reward function; $H$ is the horizon that indicates the end of trajectories at the $H$-th step; $\gamma$ is the discount factor. The entire decision trajectory can be denoted with a sequence of transition tuples $\tau=\{(s_{t},a_{t},r_{t})\}_{t=0}^{H-1}$, where $(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$, $r_{t}\sim R(s_{t},a_{t})$ and $s_{t+1}\sim T(s_{t},a_{t})$. Nevertheless, trajectory modeling defined under these Markovian assumptions may result in limited performance as ablated in Section 5.6, as it enforces temporally causal relations within trajectories, potentially sacrificing global consistency.
Source and Target Domains.
Adapting the standard MDP with domain-dependent annotations as $\mathcal{M}(\Omega):=\langle\mathcal{S}_{\Omega},\mathcal{A}_{\Omega},T_{\Omega},R_{\Omega},H,\gamma\rangle$ (Niu et al., 2024), where $\mathcal{S}_{\Omega}$, $\mathcal{A}_{\Omega}$, and $T_{\Omega}$ represent domain-dependent state spaces, action spaces, and transition dynamics, respectively. The reward function $R_{\Omega}$, influenced by $\mathcal{S}_{\Omega}$ and $\mathcal{A}_{\Omega}$, is also task-relevant. Here, $\Omega$ denotes the domain, which is shaped by the environment and the embodiments. Source domain(s) $\Omega_{src}=\{\Omega_{src}^{i}\}^{N}_{i=1}$, $N\in\mathbb{N}^{+}$, and target domain $\Omega_{tgt}$ may differ in state, action, and reward spaces, as well as in transition dynamics, reflecting the complex interplay of visual/viewpoint and dynamics gaps in the real world.
Diffusion Models.
Diffusion models have emerged as a powerful class of generative models (Ho et al., 2020), which typically involves solving two stochastic differential equation (SDE) phases: the forward and the reverse processes. The forward process models the gradual addition of noise to the data, transforming the data distribution $p(\mathbf{x}_{0})$ into a noised distribution over a series of discrete time steps $K$. This process can be described by the following equation:
$q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}(\mathbf{x}_{k};\sqrt{\alpha_{k-1}}\mathbf{x}_{k-1},(1-\alpha_{k-1})\mathbf{I}),\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),k\in[1,\cdots,K]$ (1)
where $\alpha_{k}$ are the variance schedules. Given $\mathbf{x}_{0}$, we could sample the results of forward SDE at any noise step $k$:
$\mathbf{x}_{k}\sim\mathcal{N}(\mathbf{x}_{k};\mu_{k},\sigma_{k}^{2}\mathbf{I}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\mathbf{x}_{0},(1-\alpha_{1}\alpha_{2}\cdots\alpha_{k})\mathbf{I})$ (2)
The reverse process, on the other hand, involves learning to denoise the data, effectively reversing the forward process. It can be modeled as:
@@
where $\mu_{\theta}(\mathbf{x}_{k},k)$ is the learnable mean of reverse conditional distribution. The training process involves optimizing the variational lower bound of the data likelihood, a process which encourages the model to accurately reconstruct the data from noise. This can boil down to simply minimizing the difference of the forward posterior and reverse conditional distribution $D_{KL}(q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})|p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}))$ across all denoising steps, where $q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})=\mathcal{N}(\mathbf{x}_{k-1};\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0}),\Sigma_{q})$. As the variance is only dependent on pre-defined $\alpha$ coefficients, we can construct the variance of the approximate reverse distribution $\Sigma_{k}$ also as $\Sigma_{q}=\sigma_{q}^{2}\mathbf{I}$. Thus, $\mu_{\theta}(\mathbf{x}_{k},k)$ is to clearly approach the forward process posterior mean $\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0})$:
$\arg\min_{\theta}D_{KL}(q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})|p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}))\Leftrightarrow\arg\min\frac{1}{2\sigma_{q}^{2}}\|\mu_{\theta}(\mathbf{x}_{k},k),\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0})\|_{2}^{2}$ (4)
In practice, we adopt a further simplified surrogate loss instead of the distribution mean matching:
@@
where $\epsilon_{\theta}$ represents a parameterized network predicting the noise based on $\mathbf{x}_{k}$ and $k$. This remains identical to estimating the forward posterior mean, by adopting necessary re-parameterization (Ho et al., 2020). In our context, the data samples $\mathbf{x}$ are assigned with decision trajectories $\tau$.
4 xTED: A Cross-Domain Trajectory Editing Framework
Figure 2: The model architecture is delicately designed for capturing intricate dependencies among transitions and decision elements (states, actions and rewards).
In Section 4.1, we introduce the design philosophies behind model architecture and analyze why it is adept at handling the intricate heterogeneity, dependencies and temporal consistency within trajectories than naïve implementations, as illustrated in Fig. 2. Next, we formalize how it effectively scaffolds the cross-domain trajectory editing paradigm, as detailed in Section 4.2.

### v3

@@ Abstract
Reusing pre-collected data from different domains is an appealing solution for decision-making tasks, especially when data in the target domain are limited. Existing cross-domain policy transfer methods mostly aim at learning domain correspondences or corrections to facilitate policy learning, such as learning task/domain-specific discriminators, representations, or policies. This design philosophy often results in heavy model architectures or task/domain-specific modeling, lacking flexibility. This reality makes us wonder: can we directly bridge the domain gaps universally at the data level, instead of relying on complex downstream cross-domain policy transfer procedures? In this study, we propose the Cross-Domain Trajectory EDiting (xTED) framework that employs a specially designed diffusion model for cross-domain trajectory adaptation. Our proposed model architecture effectively captures the intricate dependencies among states, actions, and rewards, as well as the dynamics patterns within target data. Edited by adding noises and denoising with the pre-trained diffusion model, source domain trajectories can be transformed to align with target domain properties while preserving original semantic information. This process effectively corrects underlying domain gaps, enhancing state realism and dynamics reliability in source data, and allowing flexible integration with various single-domain and cross-domain downstream policy learning methods. Despite its simplicity, xTED demonstrates superior performance in extensive simulation and real-robot experiments.
Machine Learning, ICML
@@ 1 Introduction
Solving real-world tasks with reinforcement learning (RL) or imitation learning (IL) often faces serious data scarcity issues (Guiochet et al., 2017; Zhan et al., 2022). To ensure reasonable performance, researchers often have to resort to laborious and costly data collection, tedious reset operation, and troublesome reward specifications (Zhu et al., 2020). Faced with restricted data acquisition, an alternative direction is to incorporate data generated with simulation or pre-collected from other domains (i.e. source domains) into policy learning for greater synergy (Niu et al., 2022, 2023; Open X-Embodiment et al., 2023). However, source domains inevitably bear some domain gaps, such as appearance gaps (Tobin et al., 2017), dynamics gaps (Peng et al., 2018), and morphology gaps for embodied agents (Gupta et al., 2021). These domain gaps greatly restrict the usability of source domain data, as direct incorporation may negatively impact policy learning (Niu et al., 2024).
@@
Existing cross-domain policy learning methods tend to design domain-specific policy transfer models equipped with domain correspondences (Zhang et al., 2020), corrections (Eysenbach et al., 2021), or discriminations (Stadie et al., 2016; Sharma et al., 2019) that either rely on task-specific architecture designs (e.g. special encoders (Mueller et al., 2018; Wang et al., 2022a) and domain-specific regularizations (Desai et al., 2020; Eysenbach et al., 2021; Niu et al., 2022; Xue et al., 2023)), or are only applicable to a specific domain of data (e.g. adaptable vision encoder design only applicable to image inputs (Rao et al., 2020). Additionally, the domain-specific discriminators or mappings (Stadie et al., 2016; Liu et al., 2018) sometimes require tedious efforts to finetune for accommodating multiple source domains, substantially hindering data reuse efficiency and efficacy. The fundamental limitation of these approaches lies in their focus on bridging domain gaps within the policy learning process rather than addressing the root cause of the problem: the domain gaps in the data itself. While existing methods attempt to compensate for these gaps through complex model architectures and additional training objectives, they inevitably introduce new complexities and restrictions. This observation motivates us to rethink: instead of adapting the policy learning process to accommodate domain gaps, can we directly bridge these gaps at the data level?
Figure 1: While sharing conceptual similarities with image editing, trajectory editing introduces distinct challenges due to the inherent complexity of sequential decision-making data, such as heterogeneous elements and complex internal dependencies .
@@
Indeed, if source data can be directly transformed to minimize domain gaps with the target domain, we can flexibly choose any downstream policy learning method based primarily on task-specific considerations, bypassing the burden of cross-domain design complexities. However, this transformation requires a careful balance: we must modify the source data to align with target domain properties while preserving its essential primitive task information that can supplement policy learning from solely using target data. Traditional data augmentation and adaptation methods are insufficient for this purpose, since they typically overlook the benefits of the source data, and that is where “editing” becomes crucial. These pursuits naturally remind us of the diffusion-based image editing (Meng et al., 2022; Kawar et al., 2023), that transforms the visual style and aesthetic elements of an image without altering its narrative content, such as converting a stroke-painted image into a photorealistic one. However, as depicted in Fig. 1, editing decision-making trajectory data from the source domain can be substantially more challenging, due to potential perception gaps and altered viewpoints in the state inputs, as well as the inherent dynamics gaps. Moreover, trajectory data in decision-making tasks are composed of fundamentally heterogeneous elements—such as observational states, actions, and rewards—that, if naïvely treated as equivalent pixels in an image, would obscure the recognition of their respective inherence and delicate internal dependencies. Thus, despite being conceptually similar, image editing methods are not directly applicable to cross-domain trajectory editing due to the intrinsic heterogeneity and dependencies within trajectories and the complex domain discrepancies.
@@
To address these challenges, we propose the Cross-Domain Trajectory EDiting (xTED) paradigm with a novel diffusion model architecture tailored for decision-making data, which captures the target trajectory distribution as a prior. The proposed architecture features separate encoding and decoding of states, actions, and rewards to preserve their distinct physical meanings and internal temporal dependencies. To effectively capture the dynamics patterns within trajectories, we introduce dependency structure modeling mechanisms that capture the internal dynamics relationships among states, actions, and rewards. With the assistance of the above designs, we successfully extend the philosophy in diffusion-based image editing (Huang et al., 2024) to decision-making data, equipping the xTED pipeline with three simple steps: (1) train diffusion model on the target data; (2) perturb source data with noises and then denoise them with the pre-trained diffusion model; (3) incorporate edited source data into target data for policy learning with any algorithm at will. Through extensive experiments, we show that incorporating source data edited with xTED consistently yields performance gains over training solely on target data, while directly adding unprocessed source data often results in significant performance degradation, particularly in real-robot manipulation tasks. As a data adaptation method, xTED enables various choices of downstream single-domain policy learning methods to address cross-domain challenges, and can also be combined with other cross-domain methods to further enhance their performance.
@@ 2.1 Cross-Domain Policy Adaptation
Addressing the challenges posed by domain gaps has long been identified as an essential task for real-world policy learning (Niu et al., 2024). The most straightforward approach is to construct direct mappings for state and action space between source and target domains (Liu et al., 2018; Kim et al., 2020; Zhang et al., 2020; Raychaudhuri et al., 2021; Wang et al., 2022b). An alternative is to learn domain-agnostic task-relevant embedding with mutual information criterion (Franzmeyer et al., 2022) and explicit domain discrimination (Stadie et al., 2016; Sharma et al., 2019). It can also be achieved with temporal contrastive learning methods (Sermanet et al., 2018; Dwibedi et al., 2019; Yang et al., 2023; Choi et al., 2023; Li et al., 2024) to learn representations that dependent on task progression while neglecting domain-variant information. Other than representation learning, it is also preferable to directly regularize the policy learning process under mild observation gaps, by reward augmentation (Eysenbach et al., 2021; Liu et al., 2022; Xue et al., 2023) or reweighting the value update (Niu et al., 2022, 2023; Xu et al., 2023a). All existing works suffer from complicated domain/task-specific designs, which lack flexibility and hinder convenient reuse and fine-tuning of those models to accommodate data from multiple source domains. Until now, there is no generic and flexible approach in the literature that handles domain gaps at the data level, and could potentially bypass the aforementioned drawbacks.
@@ 2.2 Diffusion Models for Decision Making
Recently, diffusion models have been adopted in various decision-making problem settings, including generating multi-modal policies (Chi et al., 2023; Wang et al., 2023), single-step transitions (Lu et al., 2023), subgoals (Black et al., 2024), trajectories (Janner et al., 2022; Ajay et al., 2023a; He et al., 2023; Carvalho et al., 2023; Luo et al., 2024), and videos for planning (Ajay et al., 2023b; Yang et al., 2024; Du et al., 2023, 2024). However, prior works primarily focus on single-domain generation tasks, such as data augmentation, rather than addressing cross-domain challenges. In contrast, xTED introduces a distinct setting by employing diffusion models to transform source trajectories to more closely resemble target domain characteristics while preserving useful primitive task information from source domains. This setting is less data-intensive, as it does not require extensive training data as needed in diffusion-based generation tasks that aim to generate data from pure Gaussian noises.
@@ 3 Preliminaries
Notations. We introduce the relevant notations and elements of decision trajectories with the standard formulation of Markov Decision Process (MDP). A finite-horizon MDP is defined as a tuple $\mathcal{M}:=\langle\mathcal{S},\mathcal{A},T,R,H,\gamma\rangle$ where $\mathcal{S}$ and $\mathcal{A}$ are state and action spaces, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition dynamics; $R:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathbb{R}}$ is the reward function; $H$ is the trajectory horizon; $\gamma$ is the discount factor. The entire decision trajectory can be denoted with a sequence of transition tuples $\tau=\{(s_{t},a_{t},r_{t})\}_{t=0}^{H-1}$, where $(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$, $r_{t}\sim R(s_{t},a_{t})$ and $s_{t+1}\sim T(s_{t},a_{t})$.
In this work, we focus on the cross-domain setting, where we could have one or many source domains, which differ from the target domain in terms of state (visual/viewpoint difference) and reward spaces as well as transition dynamics. Our goal is to adapt these source domain data to align with the target domain where preserving their primitive task information to facilitate target domain policy learning.
Diffusion Model. Diffusion models have emerged as a powerful class of generative models (Ho et al., 2020), which typically involve two phases: the forward and the reverse processes. The forward process models the gradual addition of noise to the data, transforming the data distribution $p(\mathbf{x}_{0})$ into a noised distribution over a series of discrete time steps $K$. This process can be described by the following equation:
$q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=$ $\mathcal{N}(\mathbf{x}_{k};\sqrt{\alpha_{k-1}}\mathbf{x}_{k-1},(1-\alpha_{k-1})\mathbf{I}),$ (1)
$\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),k\in[1,\cdots,K]$
where $\alpha_{k}$ are the variance schedules. Given $\mathbf{x}_{0}$, we could sample the results of forward process at any noise step $k$:
$\mathbf{x}_{k}$ $\sim\mathcal{N}(\mathbf{x}_{k};\mu_{k},\sigma_{k}^{2}\mathbf{I})$ (2)
$:=\mathcal{N}(\mathbf{x}_{k};\sqrt{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\mathbf{x}_{0},(1-\alpha_{1}\alpha_{2}\cdots\alpha_{k})\mathbf{I})$
The reverse process, on the other hand, involves learning to denoise the data, effectively reversing the forward process:
@@
where $\mu_{\theta}(\mathbf{x}_{k},k)$ is the learnable mean of reverse conditional distribution. The training process involves optimizing the variational lower bound of the data likelihood, a process that encourages the model to accurately reconstruct the data from noise. This can boil down to minimizing the difference of the forward posterior and reverse conditional distribution $D_{KL}(q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})|p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}))$ across all denoising steps, where $q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})=\mathcal{N}(\mathbf{x}_{k-1};\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0}),\Sigma_{q})$. As the variance is only dependent on pre-defined $\alpha$ coefficients, we can construct the variance of the approximate reverse distribution $\Sigma_{k}$ also as $\Sigma_{q}=\sigma_{q}^{2}\mathbf{I}$. Thus, $\mu_{\theta}(\mathbf{x}_{k},k)$ can be optimized towards the forward process posterior mean $\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0})$ as:
$\arg\min_{\theta}D_{KL}(q(\mathbf{x}_{k-1}|\mathbf{x}_{k},\textbf{x}_{0})|p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}))$ (4)
$\Leftrightarrow\arg\min\frac{1}{2\sigma_{q}^{2}}\|\mu_{\theta}(\mathbf{x}_{k},k)-\mu_{q}(\mathbf{x}_{k},\mathbf{x}_{0})\|_{2}^{2}$
In practice, we can adopt a further simplified surrogate loss instead of the distribution mean matching (Ho et al., 2020):
@@
where $\epsilon_{\theta}$ represents a learnable network predicting the noise based on $\mathbf{x}_{k}$ and $k$, which is equivalent to estimating the forward posterior mean through re-parameterization trick.
Diffusion-based image editing. Image editing methods aim to modify visual styles while preserving the semantic content in source images, striking a balance between style realism in the target domain and content faithfulness to the source image. Unlike standard diffusion generative models that denoise fully noised inputs, diffusion editing methods (Meng et al., 2022) observe that this desired partial modifications can be achieved by initiating the reverse process from an intermediate noise step $k<K$ of forward process as in Eq. (2). This approach allows for selective adjustments in the source image, akin to how we want to transform source trajectories: removing domain biases while retaining valuable task-relevant information. However, directly applying diffusion-based editing to decision-making data faces some technical barriers: 1) trajectories contain heterogeneous elements (states/actions/rewards) with distinct physical meanings, and 2) intricate dependencies exist among those elements. These challenges demand specially designed diffusion model architectures and tailored editing mechanisms as compared to editing in the image domain.
4 xTED: Cross-Domain Trajectory Editing
Figure 2: The model architecture is designed for capturing heterogeneous physical meanings of decision-making elements (states, actions, and rewards) and their intricate temporal and internal dependencies.
In this section, we describe the detailed diffusion model architecture of xTED and its design philosophy to handle the intricate heterogeneity, sequential characteristics, and internal dependencies within the trajectory data, as illustrated in Fig. 2. Next, we describe how to use it to effectively scaffold the cross-domain trajectory editing paradigm.